OpenStudy (credmond):

Write the complex number in polar form with argument θ between 0 and 2π. 7+7(sqrt)3i

OpenStudy (anonymous):

\[7+7\sqrt{3}i\] in polar form you need two numbers, \(r=\sqrt{a^2+b^2}\) and \(\theta\) in this example you can use \(\theta=\tan^{-1}(\frac{b}{a})\) because you are in quadrant 1

OpenStudy (anonymous):

theta you can just about do in your head, it is \(\frac{\pi}{3}\) but if it is not clear we can get that second \(r\) is routine, it is \(\sqrt{7^2+7\sqrt{3})^2}=\sqrt{49+3\times 7^2}=\sqrt{4\times 7^2}=14\)

OpenStudy (anonymous):

that was king of show offy, you can also just compute \[\sqrt{7^2+(7\sqrt{3})^2}=\sqrt{49+147}=\sqrt{196}=14\]

OpenStudy (anonymous):

and \(\theta=\tan^{-1}(\frac{\sqrt{3}}{2})=\frac{\pi}{3}\)

OpenStudy (anonymous):

you got this?

OpenStudy (credmond):

AHHH thank you sooo much!! I really appreciate it!

OpenStudy (credmond):

I just got caught up on the algebra in the square root. THANK YOU!!

OpenStudy (anonymous):

yw